# Globally Rigid Convex Braced Polygons

### v1

@@ 3.4 Main theorems
The following are equivalent for a minimally 3-connected braced polygon graph $G$.
@@
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are infinitesimally rigid.
@@
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are convexly rigid.
@@
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are globally rigid.
@@
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are super stable.
@@
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are properly stressed.
@@
    $G$ has $n$ boundary vertices and $n-2$ internal braces ($n\geq 4$).
Proof. The equivalence between (a) and (b) follows from Theorem 2.1. (c) implies (b) by definition, and (d) implies (c) by Connelly’s super stability theorem. (e) implies (d) by Theorem 3.1. (g) implies (e) by Lemma 3.2. Finally, Suppose (a) holds. By Theorem 3.3, $G$ has an infinitesimally rigid strictly convex framework $(G,{\mathbf{p}})$ with a proper stress. Hence, $m\geq 2n-3+1=2n-2$. By Lemma 3.2, $G$ is a circuit and (g) follows. Similarly, Lemma 3.2 shows that (g) and (h) are equivalent. ∎
Geleji and Jordán [Geleji and Jordán(2013), Theorem 1.2] gave a criterion, known as the “unique interval property”(See Appendix), that is equivalent to all convex configurations of a polygonal circuit having proper stress, we show that minimal 3 connectedness is necessary and sufficient.
@@
Let $G$ be a braced polygonal graph with $n$ vertices and $22n-$ 2 edges. Then, $G$ is minimally 3-connected if and only if all strictly convex polygon frameworks $(G,{\mathbf{p}})$ are properly stressed.
Proof. Suppose $G$ is minimally 3-connected. By Lemma 3.2, $G$ is a circuit. By Theorem 3.6, all strictly convex polygon frameworks of $G$ are properly stressed.
@@
The condition of having $n-2$ braces guarantees the existence of a stress in any configuration. Then the minimal 3-connectivity prevents the stress on any edge from changing its sign. Does having a stress in a minimally 3 connected convex braced polygon imply it is properly stressed in that particular configuration when there are not enough braces? The answer is negative, as shown in Figure 10. This shows that it is necessary for the stress to vary continuously with respect to the configuration.
Figure 10: This braced polygon is minimally 3 connected. However, $e_{47}$ has zero stress when vertices $\{1,2,3,4,5,6\}$ are placed on a circle and vertices $\{6,7,8,9,10,1\}$ are placed on a circle. The set of convex stressed configurations is disconnected for the braced polygon graph so we cannot vary the stress continuously from a properly stressed configuration to this figure.
In order to interpolate between our result and the work of Geleji and Jordán, in the appendix we give a combinatorial proof for the equivalence between the unique interval property and minimal 3-connectivity for circuits.
Figure 11: Relation among properties of braced polygonal graphs.
You might wonder if minimal 3-connectivity is equivalent to other properties listed in Theorem 3.6 for circuits. There are non-minimally 3-connected circuits which satisfy (a), (b), (c), or (d) in Theorem 3.6. Examples are given in the next subsection. Figure 11 summarizes the results.
3.5 Braced Polygon Graph Circuits with 7 or 8 vertices
As a concrete example and application of our theorems, in this subsection we shall examine strictly convex braced polygons with $7$ vertices and $12$ edges. Figures 1214, and 15 show the examples of strictly convex braced polygons, with $7$ vertices and $12$ edges, collected into the catagories of being always super stable with a proper stress (Figure 12), always super stable, but with some configurations that have cables for internal braces (Figure 14), and those that are not globally rigid in the plane, but are convexly rigid in the plane (Figure 15).
Figure 12: These are all the $7$-convex braced circuits with a proper stress. All the internal braces are struts in all configurations. See Theorem 3.7.
Figure 13: These are all the $8$-convex braced circuits with a proper stress. All the internal braces are struts in all configurations. See Theorem 3.7.
Figure 14: These are all the graphs of braced 7-gons, where the configuration is super stable in all strictly convex realizations, but where some of the (braced) edges of the configurations are not always struts. Each red edge may be a cable instead of strut in some configurations.
Note that the red edges in Figure 14 are defined as those internal bracing edges that do not always have a negative sign for strictly convex configurations. It is necessary, in that case, that the edge does not have a degree $3$ vertex, at one end, but it is not always sufficient. Indeed, in the last two cases in the second column, there is an edge that has both vertices of degree $3$, but it is not colored red. In order for a stress on an edge to remain negative for all convex configurations, from Theorem 3.7, it is necessary and sufficient, that, when the edge is removed, the graph should fail to be $3$-connected.
Figure 15: These are all the $7$-gon braced circuits in the plane, which are convexly rigid, i.e. unique in all convex realizations, but are not globally rigid in the plane. The blue vertices may have realizations in the plane, where they are inside the convex hull of other vertices and, in that case, the configuration is not globally rigid in the plane. The idea is to move the green vertices to the interior of the corresponding red edge, and then there is a $2$-dimensional space of stresses, indicating a non-trivial infinitesimal motion, such that one of the implied equivalent but not congruent configurations is convex and the other is not. See Figure 6, which is related to the second framework here.
@@ 4 Examples and Questions
For a graph G, with a Hamiltonian cycle, that is generically rigid in the plane, is there a reasonable polynomial-time algorithm to determine when all strictly convex configurations, for the Hamiltonian cycle, are infinitesimally rigid in the plane?
Question 4.2.
Question 4.3.
@@
One idea is that two graphs known to be super stable in all convex configurations can be “stuck” together through superposition to create another super stable graph. When two graphs share 3 vertices not on a line, an edge that has a positive stress can cancel with an edge with a negative stress. This is shown in Figure 17 where a heptagon is assembled by adding three smaller polygons while canceling two edges stresses. Note that here one (but not both) of the red edges could be a cable depending on the configuration.
Figure 17: An example of assembling larger super stable graphs from smaller graphs. The cancelled edges have opposite signs in all convex configurations in the components.
@@
Sometimes, it is possible to have super stable circuit framework configurations as in Figure 18, where the vertices are in a strictly convex position, but not all the boundary edges of the corresponding convex polygon are part of the framework. Can you see how the superposition works here? In such a convex position, the cables and struts always have the corresponding positive and negative stresses, when in a strictly convex position.
@@
Figure 18: These tensegrities are always super stable, via superposition, when their vertices are in a strictly convex configuration. The top part of Figure 19 shows the derivation for one of the cases.
Figure 19: These tensegrities are always super stable, via superposition, when their vertices are in a strictly convex configuration. Here the red edges are the ones that always cancel. Notice that the tensegrity of the bottom configuration is the same as the Grünbaum tensegrity of Figure 20, except that the $84$ and $51$ edges have been switched in the plane projection.
@@
For all strictly convex braced polygons with seven or fewer vertices, we can decide whether there is an example that is not globally rigid in the plane as a bar framework, or in all cases it is super stable. However, we have found some strictly convex cases with $8$ vertices, where we know that the framework is not super stable, and by Theorem $12.1$ in [Connelly and Gortler(2015)] not even universally rigid in some strictly convex cases. But we cannot decide whether it is globally rigid or not for the non-universally rigid strictly convex cases in the plane. The graph of Figure 20 is such a case in the plane projection, when the outside polygon is strictly convex. It looks like the examples of Figure 20 are always globally rigid as bar frameworks in the plane, but we have no proof.
@@
In the case of the Grünbaum bar framework (Figure 20), the $1458$ and $2367$ vertices each form an infinitesimally rigid framework in the plane, since they are constructed of two triangles with a common edge in the plane. If there is another configuration of these (infinitesimally) rigid subsets of the configuration, then the other non edges $15$ and $37$ are implicitly cables, since they can only decrease in length. If either one of these lengths does not decrease in length, then the bottom row of Figure 19 shows that with this added bar, the configuration would be universally rigid. Nevertheless, with this added information, we still don’t see how the Grünbaum configurations are always globally rigid in the plane.
Figure 20: This shows a $3$-dimensional tensegrity that projects orthogonally into the plane. It is a superposition of four planar quadrilaterals shown on the right. In three-space it is super stable, but in the plane it is only universally rigid.
@@
Figure 21 shows that there is a region of the space of convex configurations of the Grünbaum bar framework that are not universally rigid, and other regions that are super stable, assuming a rotational symmetry of 90 degrees. The red region corresponds to the positon of vertex $7$, where the configuration is not universally rigid, but the polygon is still strictly convex. It turns out that when the point $7$ is below the $2,6$ line in Figure 21, the configuration is not universally rigid.
@@
Figure 21: These tensegrities are various configurations of the Grübaum tensegrity in the plane as in Figure 20. They are all infinitesimally rigid as bar frameworks, but those members that have 00 stress are omitted to further understand whether the stress matrix is PSD or not. But the 00 stressed bars still contribute to the bar framework’s infinitesimal rigidity. Thus, the stresses of the stress matrix, under suitable normalization, will change continuously and so will the eigenvalues of the stress matrix.
@@
A This essentially is a square with two struts as diagonals, and so it is super stable.
@@ C Point $7$ and all the other points lie on the same circle and so there is a
C reflection symmetry in the whole configuration that has proper stress with
C $\omega_{84}=\omega_{62}=0$ and all other stresses on the internal edges negative.
@@ D These configurations must be super stable until Point $7$ crosses the horizontal line,
D and $\omega_{84}=\omega_{62}>0$ by looking ahead to configurations E.
E This is the Grünbaum tensegrity in Figure 20,
E so the stress matrix $\Omega$ is PSD with only one extra 00 eigen value.
@@ F The circular arc of the red region is part of a circle with center at the midpoint
F between the points $6$ and $8$. So the angle at Point $7$ is at most a right angle and the
F configuration is convex. Looking ahead to the configuration I, the 00 eigen value became negative.
@@ I $\omega_{67}=0$ and so the value of the stress energy is the negative of the stress
I energy of a square with diagonals as struts.
Cases from Figure 21
Remark 4.1.
Another simple way to create examples of universally rigid frameworks is to adjoin a sequence of universally rigid “plates” together with a strut from the first to last as in Figure 22.
Figure 22: The tensegrities on the left, thought of as bar frameworks, are made of rigid triangles placed edge-to-edge with a strut connecting the first to last going through all the interior of the triangles, and thus easily seen to be universally rigid, just by the triangle inequality. This shows how to construct such a framework as a superposition of simpler tensegrities. The red and green members, when superimposed, have their stresses cancel.
One of the problems with applying this idea more generally is that each of the plates needs to be universally rigid, at least as a bar framework. If there is an open set of configurations that is universally rigid, except for a triangle in the plane, if more than one of the plates is not a triangle, it will have a non-zero stress, and then the union will have a $2$-dimensional set of stresses. We are treating the case of circuits, here, which have only a one-dimensional space of stresses. The last case here in Figure 22 is the next to last example in the first column, in the table of Figure 14.

### v2

@@ 3.4 Main theorems
Let $n\geq 4$. The following are equivalent for a minimally 3-connected braced polygon graph $G$ with $n$ vertices.
@@
    All strictly convex braced polygons $(G,{\mathbf{p}})$ are infinitesimally rigid.
@@
    All strictly convex braced polygons $(G,{\mathbf{p}})$ are convexly rigid.
@@
    All strictly convex braced polygons $(G,{\mathbf{p}})$ are globally rigid.
@@
    All strictly convex braced polygons $(G,{\mathbf{p}})$ are super stable.
@@
    All strictly convex braced polygons $(G,{\mathbf{p}})$ are properly stressed.
@@
    $G$ has $n-2$ internal braces.
Proof. The equivalence between (a) and (b) follows from Theorem 2.1. (c) implies (b) by definition, and (d) implies (c) by Connelly’s super stability theorem. (e) implies (d) by Theorem 3.1. (g) implies (e) by Lemma 3.2. Finally, Suppose (a) holds. By Theorem 3.3, $G$ has an infinitesimally rigid strictly convex framework $(G,{\mathbf{p}})$ with a proper stress. Hence, $m\geq 2n-3+1=2n-2$. By Lemma 3.2, $G$ is a generic rigidity circuit and (g) follows. Similarly, Lemma 3.2 shows that (g) and (h) are equivalent. ∎
Geleji and Jordán [Geleji and Jordán(2013), Theorem 1.2] gave a criterion, known as the “unique interval property”(See Appendix), that is equivalent to all convex configurations of a braced polygonal circuit having a proper stress. We show that this is also equivalent to minimal 3-connectedness.
@@
Let $G$ be a braced polygon graph with $n$ vertices and $22n-$ 2 edges. Then, $G$ is minimally 3-connected if and only if all strictly convex braced polygons $(G,{\mathbf{p}})$ are properly stressed.
Proof. Suppose $G$ is minimally 3-connected. By Lemma 3.2, $G$ is a generic rigidity circuit. By Theorem 3.6, all strictly convex braced polygon of $G$ are properly stressed.
@@
The condition of having $n-2$ braces guarantees the existence of a stress in any strictly convex configuration. Then the minimal 3-connectivity prevents the stress on any edge from changing its sign. Does having a stress in a minimally 3-connected convex braced polygon imply the existance of a proper stress in that particular configuration when there are not enough braces? The answer is negative, as shown in Figure 10. This shows that it is necessary for the stress to vary continuously with respect to the configuration.
Figure 10: The underlying braced polygon graph of this framework is minimally 3-connected. However, $e_{47}$ has zero stress when vertices $\{1,2,3,4,5,6\}$ are placed on one circle and vertices $\{6,7,8,9,10,1\}$ are placed on another circle. The set of convex stressed configurations is disconnected for this braced polygon graph so we cannot vary the stress continuously from a properly stressed configuration to this configuration.
In order to interpolate between our result and the work of Geleji and Jordán, we will give a combinatorial proof for the equivalence between the unique interval property and minimal 3-connectivity for generic rigidity circuits in the appendix.
Figure 11: Relation among properties of braced polygon graphs.
The reader may wonder if all braced polygon graphs which satisfy one or more of the properties listed in Theorem 3.6 are minimally 3-connected. This is not the case. There are generic rigidity circuits which satisfy (a), (b), (c), or (d) in Theorem 3.6 that are not minimally connected. Figure 11 summarizes the results.
@@ 4 Examples and Questions
In the figures of braced polygonal graphs throughout this section, we will adopt the convention that all strictly convex plane realisations of the graphs shown have a stress which is positive on the dashed blue edges and negative on the solid blue edges. These stresses take positive values on red edges at some configurations and negative values on red edges at other configurations.
4.1 Superposition technique
Consider the following basic question for convex braced polygons.
@@ 4.1 Superposition technique
One idea is that two braced polygon graphs known to be super stable in all convex configurations can be “stuck” together through superposition to create another super stable braced polygon graph. When two such graphs share 3 vertices not on a line, an edge on the bounding polygon of one (that necessarily has a positive stress) can cancel with a brace in the other (that necessarily has a negative stress). This is shown in Figure 12 where a convex braced heptagon graph is assembled by adding three smaller braced polygon graphs while deleting two edges. The fact that the resulting braced polygon graph is super stable in all convex configurations follows from the following result of Connelly, which can be proved using the same techniques as his proof of [Connelly(1982), Lemma 4].
Lemma 4.1.
Let $(G,p)$ be a framework in general position in $\mathbb{R}^{d}$ and $G_{1},G_{2}$ be subgraphs of $G$ such that $G=G_{1}\cup G_{2}$ and $|V(G_{1})\cap V(G_{2})|\geq d+1$. Suppose that $G_{i}$ is super stable for some equilibrium stress $\omega_{i}$ for both $i=1,2$. Let $\omega:E(G)\to\mathbb{R}$ by putting $\omega(e)=\omega_{1}(e)+\omega_{2}(e)$, and taking $\omega_{i}(e)=0$ when $e\not\in E(G_{i})$. Then $\omega$ is an equilibrium stress for $G$ and $G$ is super stable for $\omega$.
Figure 12: An example of assembling a braced polygon graph which is super stable in all strictly convex configurations from three smaller ones.
@@
The technique is not restricted to convex polygons, and sometimes it is possible to have super stable frameworks as in Figure 13, where the vertices are in a strictly convex position, but not all the boundary edges of the corresponding convex polygon are part of the framework. Can you see how the superposition works here? In such a convex position, the boundary polygon edges and braces always have the corresponding positive and negative stresses, when in a strictly convex position.
@@
Figure 13: These frameworks are always super stable, via superposition, when their vertices are in a strictly convex configuration. The top part of Figure 19 shows the derivation for one of these frameworks.
4.2 Braced polygonal circuits with 7 vertices
As a concrete example and application of our observations, we examine all braced polygonal circuits with $7$ vertices. Figures 1415, and 16 show the examples of all braced polygonal circuits with $7$ vertices, collected into the categories of being always super stable with a proper stress (Figure 14), always super stable, but with some configurations that have positive stresses for internal braces (Figure 15), and those that are not globally rigid in the plane, but are convexly rigid in the plane (Figure 16). The lists are constructed based on Theorem 3.6, Theorem 3.7, and the superposition technique.
Figure 14: These are all the braced heptagonal circuits with a proper stress in all strictly convex configurations. See Theorem 3.7.
Figure 15: These are all the braced heptagonal circuits which are super stable in all strictly convex configurations, but where some of the braces of the configurations may have a positive stress.
Figure 16: These are all the braced heptagonal circuits, which are convexly rigid but are not globally rigid at all strictly convex plane configurations. The blue vertices may have realizations in the plane, where they are inside the convex hull of other vertices and, in that case, the configuration is not globally rigid in the plane. The idea is to move the green vertices to the interior of the corresponding red edge, and then there is a $2$-dimensional space of stresses, indicating a non-trivial infinitesimal motion, such that one of the implied equivalent but not congruent configurations is convex and the other is not. See Figure 6, which is related to the second framework here.
The red edges in Figure 15 are defined as those internal braces that do not a negative stress in all strictly convex configurations. It is necessary, in that case, that both end-vertices of the edge have degree at least four but this is not always sufficient. Indeed, in the last two cases in the second column, there is an edge that has both end-vertices of degree four which is not colored red. In order for a stress on an edge to remain negative for all convex configurations, from Theorem 3.3, it is necessary and sufficient, that, when the edge is removed, the graph should fail to be $3$-connected.
4.3 Braced polygonal circuits with 8 vertices
@@
In Figure 17, we give a list of braced polygonal circuits with 8 vertices satisfying the condition of Theorem 3.7. For all strictly convex braced polygons with seven or fewer vertices, we can decide whether there is an example that is not globally rigid in the plane as a bar framework, or in all cases it is super stable. However, we have found some strictly convex cases with $8$ vertices, where we know that the framework is not super stable, and by Theorem $12.1$ in [Connelly and Gortler(2015)] not even universally rigid in some strictly convex cases. But we cannot decide whether they are globally rigid or not for the non-universally rigid strictly convex cases in the plane. A notable example is the underlying graph of Figure 18, which is the orthogonal projection of a 3-dimensional framework given by Grünbaum [Grünbaum and Shephard(1975)] onto the plane. It appears that all the strictly convex plane configurations of this graph are globally rigid, but we have no proof.
Figure 17: These are all the braced polygonal circuits with 8 vertices having a proper stress in all strictly convex configurations. See Theorem 3.7.
@@
In this graph, the vertex sets $\{1,4,5,8\}$ and $\{2,3,6,7\}$ each induces an infinitesimally rigid framework in any strictly convex configuration in the plane, since they are constructed of two triangles with a common edge. If there is another configuration of these (infinitesimally) rigid subsets of the configuration, then the other non edges $15$ and $37$ can only decrease in length. If either one of these lengths does not decrease in length, then the bottom row of Figure 19 shows that with this added bar, the graph would be universally rigid at all strictly convex configurations. Nevertheless, with this added information, we still don’t see how to decide whether all the strictly convex configurations of the original graph are globally rigid.
Figure 18: The left framework is the orthogonal projection of a $3$-dimensional framework from Grünbaum and Shephard’s “Lectures in Lost Mathematics” [Grünbaum and Shephard(1975)] onto the plane. It has a 4-fold rotational symmetry and edge 78 is parallel to edge 26. It is a superposition of the four planar quadrilaterals shown on the right. The original 3-dimensional framework is known to be super stable, see [Connelly and Guest(2022), Section 11.7], but its projection onto the plane is only universally rigid. We do not know whether all the strictly convex plane realisations of its underlying graph are globally rigid.
Figure 19: These graphs are always super stable, via superposition, when their vertices are in a strictly convex configuration. Here the red edges are the ones that always cancel. Notice that the bottom graph is the same as the graph of the Grünbaum framework of Figure 18, except that the $84$ and $51$ edges have been switched in the plane projection.
@@
Figure 20 shows that there is a region of the space of convex configurations of the graph of the Grünbaum framework that are not universally rigid, and other regions that are super stable, assuming a rotational symmetry of 90 degrees. The red region corresponds to the positon of vertex $7$, where the configuration is not universally rigid, but the polygon is still strictly convex. It turns out that when the point $7$ is below the $2,6$ line in Figure 20, the configuration is not universally rigid.
@@
Figure 20: These frameworks are various plane configurations of the braced polygon graph given in Figure 18. The edges that have 00 stress are omitted. The 00 stressed bars still contribute to the bar framework’s infinitesimal rigidity, and they are all infinitesimally rigid as bar frameworks if we restore the omitted edges. See Table 1 for more detailed analysis of each case.
@@
A This essentially is a square with two braces as diagonals, and so it is super stable.
@@ C Point $7$ and all the other points lie on the same circle and so there is a
reflection symmetry in the whole configuration that has proper stress with
$\omega_{84}=\omega_{62}=0$ and all other stresses on the internal edges negative.
@@ D These configurations must be super stable until Point $7$ crosses the horizontal line,
and $\omega_{84}=\omega_{62}>0$ by looking ahead to configurations E.
E This is the Grünbaum framework in Figure 18,
so the stress matrix $\Omega$ is PSD with only one extra 00 eigenvalue.
@@ F The circular arc of the red region is part of a circle with center at the midpoint
between the points $6$ and $8$. So the angle at Point $7$ is at most a right angle and the
configuration is convex. Looking ahead to the configuration I, the 00 eigenvalue became negative.
@@ I $\omega_{67}=0$ and so the value of the stress energy is the negative of the stress
energy of a square with diagonals as braces.
Table 1: Case Analysis from Figure 20.
Another simple way to create examples of universally rigid frameworks is to adjoin a sequence of universally rigid “plates” together with a brace from the first to last as in Figure 21. One of the problems with applying this idea more generally is that each of the plates needs to be universally rigid, at least as a bar framework. If there is an open set of configurations that is universally rigid, except for a triangle in the plane, if more than one of the plates is not a triangle, it will have a non-zero stress, and then the union will have a $2$-dimensional set of stresses. We are treating the case of circuits, here, which have only a one-dimensional space of stresses. The last case here in Figure 21 is the next to last example in the first column, in the table of Figure 15.
Figure 21: The frameworks on the left are made of rigid triangles placed edge-to-edge with a brace connecting the first to last going through the interior of all the triangles, and thus are easily seen to be universally rigid, just by the triangle inequality. This shows how to construct such a framework as a superposition of simpler frameworks. The red and green members, when superimposed, have their stresses cancel.
Question 4.2.
Is there a strictly convex configuration of the braced polygon graph in Figure 18 that is not globally rigid?
Note that this graph is infinitesimally rigid at all strictly convex configurations and thus the stress matrix (and its eigenvalues) vary continuously as the configuration varies continuously in the plane. Using this fact, we can look for non-universally rigid configurations in the plane close to that given in Figure 18. The configuration in Figure 18 is universally rigid (this can be proved using a careful analysis of the superposition technique), but not super stable (because there is a 3-dimensional framework having the same one-dimensional space of stresses). However, there are small perturbations of the configuration in the plane, where the stress matrix picks up a negative eigenvalue, and as a result, we loose universal rigidity. A random search among strictly convex configurations for this graph in the plane is unlikely to find these non-universally rigid configurations, even though they form an open subset of such configurations.
4.4 Further questions for braced polygonal circuits
The central question, possibly wishful thinking, is whether there is a way to determine global rigidity for braced polygonal circuits, by the methods presented here. We have a way to tell when such a circuit is NOT globally rigid by Theorem 2.1, such as the example in Figure 6. Minimally 3-connected cases can be solved using Theorem 3.7, such as those in Figure 14. When neither of the theorems apply, our best hope is to use the superposition technique in Figure 12. This technique works for most of the cases that we have tried. There are still open cases such as Figure 18. This leads to the following:
Question 4.3.
If a braced polygonal circuit $G$ has a strictly convex plane realisation $(G,{\bf p})$ which is not globally rigid, is it true that there is a convex, but possibly not strictly convex, realisation $(G,{\bf q})$ with a non-trivial infinitesimal flex ${\bf q}^{\prime}$ such that ${\bf p}={\bf q}+{\bf q}^{\prime}$?
We are asking if every non globally rigid, strictly convex plane realisation of a braced polygonal circuit can be constructed as in Figures 1 and 6. If this is true, then the braced octagonal graph of Grünbaum, Figure 18 would be globally rigid at all strictly convex configurations.